# Unsupervised Model Learning for Quantum Impurity Systems

### v1

@@ Acknowledgements.
Appendix A Convexity of the loss function
@@
$=\mathcal{Z}_{\rm{bath}}\det_{ij}\left[V^{\sigma_{i}}_{k_{i}}V^{\sigma^{\prime}_{j}*}_{k^{\prime}_{j}}\langle T_{\tau}\hat{c}^{\dagger}_{k_{i}\sigma_{i}}(\tau_{i})\hat{c}_{k^{\prime}_{j}\sigma^{\prime}_{j}}(\tau^{\prime}_{j})\rangle_{\rm{bath}}\right]$
$=\mathcal{Z}_{\rm{bath}}\det(\Delta^{(x)})\;,$ (72)
@@
$\mathcal{Z}/\mathcal{Z}_{\rm{bath}}$ $=\sum_{x}\sum_{\{\alpha\}}\Lambda_{\{\alpha\}_{x}}\det(\Delta_{x})e^{-\langle\hat{H}^{\rm{imp}}\rangle_{\{\alpha\}_{x}}}\;,$ (75)
@@
and $\langle\hat{H}_{\rm{imp}}\rangle_{\{\alpha\}_{x}}$ the average impurity energy over a diagram
@@
$\mathcal{Z}=\mathcal{Z}_{\rm bath}\sum_{x}\sum_{\{\alpha\}_{x}}w(\{\alpha\}_{x})\;,$ (79)
@@
$\partial_{i}\log(\mathcal{Z})$ $=\frac{1}{\mathcal{Z}}\partial_{i}\bigg{[}\mathcal{Z}_{\rm bath}\sum_{x}\sum_{\{\alpha\}_{x}}e^{-\langle\hat{H}^{\rm{imp}}\rangle_{\{\alpha\}_{x}}}\Lambda_{\{\alpha\}_{x}}\det(\Delta_{x})\bigg{]}$
$=\frac{1}{\mathcal{Z}}\bigg{[}\mathcal{Z}_{\rm bath}\sum_{x}\sum_{\{\alpha\}_{x}}\partial_{i}w(\{\alpha\}_{x})\bigg{]}\;.$ (82)
@@
$-\partial_{i}\log(\mathcal{Z})$ $=-\frac{1}{\mathcal{Z}}\bigg{[}\mathcal{Z}_{\rm bath}\sum_{x}\sum_{\{\alpha\}_{x}}\partial_{i}w(\{\alpha\}_{x})\bigg{]}$ (90)
@@
Appendix B Double extended basis
To construct the projectors of the 2e-imp effective Hamiltonian, we require the basis |$\ket{Q,S,T,s_{z};m}$ on the double extended Fock space $\mathcal{H}^{\rm bath}_{1}\otimes\mathcal{H}^{\rm bath}_{0}\otimes\mathcal{H}^{\rm imp}_{\rm eff}$. The first step in the construction is to determine the charge $\hat{Q}$ and total spin $\operatorname{\hat{\mathbf{S}}}^{2}$ eigenstates |$\ket{Q,S}$. For the occupation basis we use the labelling convention:
@@ Appendix B Double extended basis
In the |$\ket{-1(1),1}$ and |$\ket{-1(1),0}$ sub-spaces we can find a common eigenbasis |$\ket{Q,S,T,s_{z};m}$ of $\hat{T}$ and $\operatorname{\hat{\mathbf{S}}}^{2}$ that allows to completely lift the multiplet degeneracy:
@@
Here we have used that
$\hat{T}\ket{0,\operatorname{\downarrow}}=-\ket{\operatorname{\downarrow},0}\;.$
Thus, the symmetric combination of |$\ket{-1(1),1(0),s_{z};m}$ states is associated to the negative $T=-1$ eigenvalue of $\hat{T}$ and the asymmetric combination to the positive eigenvalue $T=1$ of $\hat{T}$.
@@
We can combine these states such that they become eigenstates of $\hat{T}$:
@@
which has the eigenvales $W\in\{0,1\}$ in the |$\ket{0,1/2,0,s_{z};m}$ subspace. Instead of introducing a new label for only this subspace, the eigenvalues for $\hat{T}$ and $\hat{W}$ are added together in the $T$ label. This gives the states:
@@
We have thereby succeeded in reducing the dimension of each subspace spanned by these labels to $M=1$, as desired.
@@ Appendix C Details of NRG calculations
All calculations in this work were carried out using Wilson’s NRG method [75, 65] using the full density matrix approach [76, 77], which allows the precise calculation of the static observables discussed here. The impurity entropy, which is used to estimate $T_{K}$ was computed using standard thermodynamic NRG [75]. For the simulation of bare and effective models, the number of kept states is $M_{K}=2000$ and we use a Wilson chain discretization parameter of $\Lambda=2.5$ at a Wilson chain length of $N=40$. Total charge and spin projection quantum numbers are exploited.

### v2

@@ Acknowledgements.
Appendix A Hybridization expansion and convexity of the loss function
@@
$=\mathcal{Z}^{\rm{bath}}\det_{ij}\left[V^{\sigma_{i}}_{k_{i}}V^{\sigma^{\prime}_{j}*}_{k^{\prime}_{j}}\langle T_{\tau}\hat{c}^{\dagger}_{k_{i}\sigma_{i}}(\tau_{i})\hat{c}_{k^{\prime}_{j}\sigma^{\prime}_{j}}(\tau^{\prime}_{j})\rangle^{\rm{bath}}\right]$
$=\mathcal{Z}^{\rm{bath}}\det(\Delta^{(x)})\;,$ (67)
@@
$\mathcal{Z}/\mathcal{Z}^{\rm{bath}}$ $=\sum_{x}\sum_{\{\alpha\}}\Lambda_{\{\alpha\}_{x}}\det(\Delta_{x})e^{-\langle\hat{H}^{\rm{imp}}\rangle_{\{\alpha\}_{x}}}\;,$ (70)
@@
and $\langle\hat{H}^{\rm{imp}}\rangle_{\{\alpha\}_{x}}$ the average impurity energy over a diagram
@@
$\mathcal{Z}=\mathcal{Z}^{\rm bath}\sum_{x}\sum_{\{\alpha\}_{x}}w(\{\alpha\}_{x})\;,$ (74)
@@
$\partial_{i}\log(\mathcal{Z})$ $=\frac{1}{\mathcal{Z}}\partial_{i}\bigg{[}\mathcal{Z}^{\rm bath}\sum_{x}\sum_{\{\alpha\}_{x}}e^{-\langle\hat{H}^{\rm{imp}}\rangle_{\{\alpha\}_{x}}}\Lambda_{\{\alpha\}_{x}}\det(\Delta_{x})\bigg{]}$
$=\frac{1}{\mathcal{Z}}\bigg{[}\mathcal{Z}^{\rm bath}\sum_{x}\sum_{\{\alpha\}_{x}}\partial_{i}w(\{\alpha\}_{x})\bigg{]}\;.$ (77)
@@
$-\partial_{i}\log(\mathcal{Z})$ $=-\frac{1}{\mathcal{Z}}\bigg{[}\mathcal{Z}^{\rm bath}\sum_{x}\sum_{\{\alpha\}_{x}}\partial_{i}w(\{\alpha\}_{x})\bigg{]}$ (85)
@@
Appendix B Extended impurity basis
To construct the projectors of the extended MC Hamiltonian, we require the basis |$\ket{Q,S,T,S_{z};m}$ on the extended Fock space Eq. 53. The first step in the construction is to determine the charge $\hat{Q}$, total spin $\operatorname{\hat{\mathbf{S}}}^{2}$ and $\hat{S}_{z}$ eigenstates |$\ket{Q,S,S_{z};m}$. For the occupation basis we use the labelling convention:
@@ Appendix B Extended impurity basis
In the |$\ket{QS}=\ket{\pm 1,1}$ and |±$\ket{\pm 1,0}$ sub-spaces we can find a common eigenbasis |$\ket{Q,S,T,S_{z};m}$ of $\hat{T}$ and $\operatorname{\hat{\mathbf{S}}}^{2}$ that allows to completely lift the multiplet degeneracy, e.g.
@@
We can combine these states such that they become eigenstates of $\hat{T}$: {adjustwidth}-1.5cm-1.5cm
@@
which has the eigenvales $W\in\{0,1\}$ in the |$\ket{0,1/2,0,S_{z};m}$ subspace. Instead of introducing a new label for only this subspace, the eigenvalues for $\hat{T}$ and $\hat{W}$ are added together in the $T$ label. This gives the states: {adjustwidth}-1.5cm-1.5cm
@@
The dimension of each subspace spanned by these labels has thereby been reduced to $M=1$, as desired.
@@ Appendix C Details of NRG calculations
All calculations in this work were carried out using Wilson’s NRG method [82, 83] using the full density matrix approach [84, 70], which allows the precise calculation of the static observables discussed here. The impurity entropy, which is used to estimate $T_{K}$ via $S_{\rm imp}(T_{K})=\tfrac{1}{2}$, was computed using standard thermodynamic NRG [82]. For the simulation of bare and effective models, the number of kept states is $M_{K}=2000$ and we use a Wilson chain discretization parameter of $\Lambda=2.5$ at a Wilson chain length of $N=40$. Total charge and spin projection quantum numbers are exploited.
Appendix D Integration with
ab initio methods
At the heart of the UML method is the computation of local observables in bare and effective models. The effective models are simple enough that this can be done accurately at any temperature using e.g. NRG. However, evaluating objects such as $\langle\hat{P}_{QS}\rangle$ in the bare model within an ab initio description is in general challenging.
For methods that operate in second quantization, we present here a straightforward way to construct the bare projectors $\hat{P}_{QS}$ that project onto the quantum number states |$\ket{Q,S}$. First we introduce the auxiliary operators,
$\hat{X}_{Q}$ $=\hat{Q}-Q$ (92)
$\hat{Y}_{S}$ $=\hat{\mathbf{S}}^{2}-S(S+1)\;,$ (93)
where $\hat{\mathbf{S}}^{2}$ is the total spin operator of the extended impurity and $\hat{Q}$ the total charge operator of the extended impurity. The operator $\hat{X}_{Q}$ eliminates all contributions with charge $Q$ to a state |$\ket{\psi}$ and similarly $\hat{Y}_{S}$ eliminates all contributions with total spin $S$. Using these operators we can now write the projector onto the multiplet subspace of $Q$ and $S$ as
where
is a normalization constant. One can similarly construct the admissibility operator $\hat{\Omega}_{\rm ad}$. However, these operators are manifestly of high order: Eq. 94 is an $N$-body operator, where $N$ is the number of particles in the extended bare impurity Hilbert space.
Within an ab initio framework, expectation values of such $N$-body operators are typically intractable, because they are usually expressed in terms of reduced density matrices (RDM) [85]. For a given basis of molecular orbitals $\{\phi_{i}\}^{M}_{i=1}$ one can express the spinless single particle RDM (1-RDM) as
$\rho_{1}(r^{\prime};r)=\sum^{M}_{i,j}~{}^{1}D^{i}_{j}\phi_{i}(r^{\prime})\phi_{j}(r)\;,$ (96)
where ${}^{1}D^{i}_{j}=d_{i}\times d_{j}=\langle\psi|\hat{c}^{\dagger}_{i}\hat{c}_{j}|\psi\rangle$ is the 1-RDM which encodes a single particle wave-function as,
$\psi(r)=\sum^{M}_{i}d_{i}\phi_{i}(r)\;.$ (97)
${}^{2}D^{pq}_{st}=2~{}^{1}D^{s}_{p}\wedge\!~{}^{1}D^{q}_{t}+~{}^{2}\Delta^{pq}_{st}\;$ (98)
where ${}^{2}\Delta$ is a cumulant that encodes the relations between multiple determinants. The tensor $D^{pq}_{st}=\langle\psi|\hat{c}^{\dagger}_{p}\hat{c}^{\dagger}_{q}\hat{c}_{s}\hat{c}_{t}|\psi\rangle$ encodes two-body correlations. The construction of the $N$-RDM is therefore possible for an arbitrary number for particles $N$, but it becomes prohibitively expensive rather quickly.
Instead of attempting to compute Eq. 94 exactly within this framework, we instead sketch a method for reducing the projectors $\hat{P}_{QS}$ to a simplified but approximate form that might be compatible with ab initio methods using 1- or 2-body RDMs.
We wish to find the best variational approximate representation to Eq. 94 using at most 2-body operators. We can do this using the FSP, Eq. 62. We therefore construct a full basis of operators for the bare Hilbert space,
$\hat{A}\in\Xi_{n}\in\{\{\hat{c}^{\dagger}_{i}\hat{c}_{j}\},~{}\{\hat{c}^{\dagger}_{i}\hat{c}^{\dagger}_{j}\hat{c}_{k}\hat{c}_{l}\},~{}\{\hat{c}^{\dagger}_{i}\hat{c}^{\dagger}_{j}\hat{c}^{\dagger}_{k}\hat{c}^{\dagger}_{l}\hat{c}_{m}\hat{c}_{n}\hat{c}_{o}\hat{c}_{p}\},\dots\}\;.$ (99)
Now retaining only the 1- and 2-body terms, we have a basis $\{\hat{h}_{i}\}$ for the approximate representation of $\hat{P}_{QS}$. For a complete basis $\Xi_{n}$, the decomposition $\hat{P}_{QS}=\sum_{i}\theta_{i}\hat{h}_{i}$ is unique and can be achieved like a regular spectral decomposition of a vector using the FSP, viz:
$\theta_{i}=\frac{\langle\hat{P}_{QS},\hat{h}_{i}\rangle}{\operatorname{|\!|}\hat{h}_{i}\operatorname{|\!|}}\;,\qquad\qquad\hat{h}_{i}\in\Xi_{n}\;.$ (100)
However, we note that the basis Eq. 99 is generally over-complete and therefore any representation of $\hat{P}_{QS}$ according to Eq. 62 is not necessarily unique. Constructing a complete operator basis on the bare Hilbert space can be challenging. On the other hand, the variational approach can find optimal representations of any operator even for an over-complete basis, and the value of the optimized loss function can serve as an error estimation for the decomposition. We argue that this is a controlled approach to obtain 2-body approximations of the UML operators $\hat{P}_{QS}$. We leave development and implementation of these methods for future work.